OpenStudy (anonymous):

dolls cost $140 per carton and trucks cost $430 per carton. If an order comes in for a total of 100 cartons for $28,500, what was the number of cartons of dolls? (hint, let T = cartons of trucks)

OpenStudy (anonymous):

Let T = cartons of trucks, Let D = cartons of dolls. From the question we know that L+T=100. And we know that D*140+T*430=28500. Solve for D or T: T=100-D. Replace that variable in the second equation: D*140+(100-D)*430=28500. 140D-430D+43000 = 28500 -290D = -14500 D = -14500/-290 D = 50. Solve for T: T = 100-D = 100-50 = 50. Plug it back into our second equation to double check: 50*140+50*430 = 28500 7000 + 21500 = 28500. Tl;dr: 50.

OpenStudy (anonymous):

eh, 2nd line should say D+T=100